A body starting from rest covers a distance of 0.45 km and acquires a velocity of 300 km/hr. Its acceleration will be? (a) 7.72 m/s^2 ans

Grade:12th pass

3 Answers

Sandeep Bhamoo
5 years ago
Apply v²-u²= 2as s= 0.45km=450m u= v=300km/hr= 250/3 m/s. On putting these values a= 625/81 m/s² please do approve my answer if correct
Khimraj
5 years ago
v = 300km/hr = 300*5/18m/s = 250/3 m/s
s = 0.45km = 450m
V2 = u2 + 2as
(250/3)2 = 0 + 2a*450 = 900a
So a = 625/81 m/s2
mehak ambreen
3 years ago
Train starts from rest with uniform acceleration of 10 cm sec-2. The velocity after one minute will be